OpenStudy (anonymous):

@austinL Solve the system of equations by elimination: {(-0.7x+0.6y=1.3), (0.5x-0.3y=-0.3)}

OpenStudy (austinl):

\(-0.7x+0.6y=1.3\) \(0.5x-0.3y=-0.3\) What I would do first is multiply the bottom equation by two. \(-0.7x+0.6y=1.3~~~~~~~~~~~~~-0.7x+0.6y=1.3\) \(2(0.5x-0.3y=-0.3)\longrightarrow\underline{~~~~~~~~~x-0.6y=-0.6}\) \(~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~0.3x=0.7\) \(x=\dfrac{7}{3}\) Then you can plug that into one of your original equations and solve for y!

OpenStudy (anonymous):

Thanks for answering. I wasn't sure if that one even posted. My computer doesn't like this site today. I actually figured it out. I meant to stay on and explain to you what I'd already done, but my computer froze again so I just gave up for a while. I was pretty sure I had an answer for this one, I just wanted to see if someone else got the same answer =) So thank you! You're always a huge help to me.